# AgentDrug: Utilizing Large Language Models in an Agentic Workflow for Zero-Shot Molecular Optimization

### v9

@@ Abstract
Molecular optimization—modifying a given molecule to improve desired properties—is a fundamental task in drug discovery. While LLMs hold the potential to solve this task using natural language to drive the optimization, straightforward prompting achieves limited accuracy. In this work, we propose AgentDrug, an agentic workflow that leverages LLMs in a structured refinement process to achieve significantly higher accuracy. AgentDrug defines a nested refinement loop: the inner loop uses feedback from cheminformatics toolkits to validate molecular structures, while the outer loop guides the LLM with generic feedback and a gradient-based objective to steer the molecule toward property improvement. We evaluate AgentDrug on benchmarks with both single- and multi-property optimization under loose and strict thresholds. Results demonstrate significant performance gains over previous methods. With Qwen-2.5-3B, AgentDrug improves accuracy by 20.7% (loose) and 16.8% (strict) on six single-property tasks, and by 7.0% and 5.3% on eight multi-property tasks. With larger model Qwen-2.5-7B, AgentDrug further improves accuracy on 6 single-property objectives by 28.9% (loose) and 29.0% (strict), and on 8 multi-property objectives by 14.9% (loose) and 13.2% (strict).
@@ 1 Introduction
[…]
@@
[…]
@@
2 Methodology
$\widehat{m}=\mathcal{M}(p||d||m)\mathrel{\hskip 3.00003pt}\text{s.t.}\mathrel{\hskip 3.00003pt}\prod_{i=1}^{N}\underset{p_{i},d_{i}}{E}(\widehat{m})=1\mathrel{\hskip 3.00003pt}\text{where}$
Here, ||||| | denotes concatenation, and $\mathbbm{1}\in{0,1}$ is an indicator function that evaluates whether the change in property $p_{i}$ satisfies the specified threshold $d_{i}$ in the intended direction.
@@ 2 Methodology
[…]
@@
[…]
@@ 3 Evaluation
[…]
@@ 4 Conclusion
This paper introduced AgentDrug, a novel approach leveraging LLMs for molecular optimization through gradient-guided refinement. By providing explicit gradient information during the refinement process, AgentDrug enables targeted molecular modifications while maintaining similarity constraints. Our experiments show that AgentDrug consistently outperforms the baselines on both single- and multi-property optimization tasks. It is particularly effective for complex multi-property objectives and can balance well with optimization accuracy and molecular similarity.
@@
Appendix A Related work
Recent advancements in AI, particularly in large language models (LLMs), have begun to transform the drug discovery process by efficiently scaling many core molecular tasks Mak and Pichika (2019). Notably, Seidl et al. (2023); Zhao et al. (2023); Liu et al. (2024b) have incrementally pre-trained LLMs to advance molecular property prediction, synthesis, and retrosynthesis. In addition, Edwards et al. (2022); Liu et al. (2023b); Li et al. (2024) have trained LLMs for molecular captioning. However, molecular optimization, which plays a critical role in designing improved compounds, remains relatively underexplored.
Table 3: Definition of the six categories of ParseError.
ParseError Definition
Appendix B Evaluation Settings
@@
Each optimization objective is expressed in natural language using a standardized prompt template (see Box A1). Specifically, for each objective, we sample 1,000 molecules from the ZINC database and train LLMs using the REINFORCE algorithm Sutton et al. (1999), where the reward function is defined as $\prod_{i=1}^{N}\underset{p_{i},d_{i}}{E}(\widehat{m})$. The models are fine-tuned using LoRA Hu et al. (2022) with a rank of 16 and a learning rate of $1\times 10^{-5}$, for a single training epoch.
Box A1: The template for wrapping an objective in natural language to form the prompt Given [a given molecule $m$], modify it to [increase or decrease] its [desired properties $\{p_{i}\}_{i=1}^{N}$] by [thresholds $\{|d_{i}|\}_{i=1}^{N}$], respectively. Importantly, the modified molecule must be similar to the given one. Respond with only the SMILES string of the modified molecule. No explanation is needed.
Appendix C Ablation studies
@@
Chat Agent Chat Agent
acc. acc. acc. acc.
@@
+LogP,+TPSA l,l 06.5 10.2 08.0 11.2
s,s 01.6 07.4 05.6 09.6
+LogP,-TPSA l,l 07.2 11.6 11.8 09.6
s,s 06.2 05.3 08.1 03.6
-LogP,+TPSA l,l 09.3 15.8 24.6 35.2
s,s 03.8 08.4 13.3 21.8
-LogP,-TPSA l,l 06.3 05.0 08.0 02.9
s,s 01.3 00.7 02.8 01.1
+LogP,+QED l,l 09.6 12.3 12.2 13.6
s,s 02.2 01.6 01.3 01.8
+LogP,-QED l,l 13.3 23.3 21.9 27.7
s,s 05.8 07.6 08.0 07.9
-LogP,+QED l,l 08.3 09.1 14.2 13.5
s,s 01.3 00.7 01.4 01.5
-LogP,-QED l,l 11.2 09.0 25.0 31.2
s,s 03.0 04.4 07.8 08.7
Average l,l 09.0 12.0 15.7 18.1
s,s 03.1 04.5 06.1 07.0
C.1 Impact of using the example molecule
We conduct an ablation study to evaluate the impact of incorporating the retrieved example molecule $m_{e}$, a component that is overlooked in prior work Liu et al. (2024a). Table 5 presents results for both AgentDrug and ChatDrug when $m_{e}$ is excluded. Even without $m_{e}$, AgentDrug/$m_{e}$ consistently outperforms ChatDrug/$m_{e}$, demonstrating the effectiveness of the remaining components in AgentDrug. Using Qwen-2.5-3B, AgentDrug/$m_{e}$ improves accuracy by 11.2% and 6.7% under loose and strict thresholds, respectively, on six single-property objectives, and by 3.0% and 1.4% on eight multi-property objectives. With Qwen-2.5-7B, improvements are 10.1% and 5.8% (single-property), and 2.4% and 0.9% (multi-property), under loose and strict thresholds, respectively.
(a) Qwen2.5-3B
(b) Qwen2.5-7B
Figure 3: Averaged results on all objectives.
C.2 Impact of the number of iterations
@@
(a) Qwen2.5-3B
(b) Qwen2.5-7B
Figure 4: Averaged results on all objectives.
We conduct an ablation study to examine how AgentDrug and ChatDrug behave under different numbers of refinement iterations $T$. Table 6 presents results for both methods using $T=4$, $5$, and $6$ iterations, respectively. The results show that AgentDrug consistently outperforms ChatDrug when using the same number of iterations. Notably, AgentDrug with fewer iterations still surpasses ChatDrug with more iterations, highlighting the superior efficiency and effectiveness of the AgentDrug framework.
@@
Appendix D Example
An illustrative example of AgentDrug’s workflow, including detailed prompts, is provided in Figure 5. The gradient signal used to guide the LLM is highlighted in red for clarity.

### v10

@@ Abstract
Molecular optimization—modifying a given molecule to improve desired properties—is a fundamental task in drug discovery. While LLMs hold the potential to solve this task using natural language to drive the optimization, straightforward prompting achieves limited accuracy. In this work, we propose AgentDrug  [^1] https://github.com/lhkhiem28/AgentDrug, an agentic workflow that leverages LLMs in a structured refinement process to achieve significantly higher accuracy. AgentDrug defines a nested refinement loop: the inner loop uses feedback from cheminformatics toolkits to validate molecular structures, while the outer loop guides the LLM with generic feedback and a gradient-based objective to steer the molecule toward property improvement. We evaluate AgentDrug on benchmarks with both single- and multi-property optimization under loose and strict thresholds. Results demonstrate significant performance gains over previous methods. With Qwen-2.5-3B, AgentDrug improves accuracy by 20.7% (loose) and 16.8% (strict) on six single-property tasks, and by 7.0% and 5.3% on eight multi-property tasks. With larger model Qwen-2.5-7B, AgentDrug further improves accuracy on 6 single-property objectives by 28.9% (loose) and 29.0% (strict), and on 8 multi-property objectives by 14.9% (loose) and 13.2% (strict).
@@ 1 Introduction
2 Methodology
[…]
Here, $||$ denotes concatenation, and $\mathbbm{1}\in{0,1}$ is an indicator function that evaluates whether the change in property $p_{i}$ satisfies the specified threshold $d_{i}$ in the intended direction.
Table 1: Definition of the six categories of ParseError.
ParseError Definition
@@ 2 Methodology
vanilla REINVENT ChatDrug $AgentDrug{\dagger}$ vanilla REINVENT ChatDrug $AgentDrug{\dagger}$
@@
vanilla REINVENT ChatDrug $AgentDrug{\dagger}$ vanilla REINVENT ChatDrug $AgentDrug{\dagger}$
@@ 3 Evaluation
ChatDrug $AgentDrug{\dagger}$ ChatDrug $AgentDrug{\dagger}$
@@
ChatDrug $AgentDrug{\dagger}$ ChatDrug $AgentDrug{\dagger}$
@@
4 Ablation Studies
4.1 Impact of using the example molecule
@@
4.2 Impact of the number of iterations
We conduct an ablation study to examine how AgentDrug and ChatDrug behave under different numbers of refinement iterations $T$. Table 6 presents results for both methods using $T=4$, 55, and $6$ iterations, respectively. The results show that AgentDrug consistently outperforms ChatDrug when using the same number of iterations. Notably, AgentDrug with fewer iterations still surpasses ChatDrug with more iterations, highlighting the superior efficiency and effectiveness of the AgentDrug framework.
(a) Qwen2.5-3B
(b) Qwen2.5-7B
Figure 3: Impact of using the example molecule. Averaged results on all objectives.
(a) Qwen2.5-3B
(b) Qwen2.5-7B
Figure 4: Impact of the number of iterations. Averaged results on all objectives.
@@ 5 Conclusion
This paper introduced AgentDrug, a novel approach leveraging LLMs for zero-shot molecular optimization through gradient-guided refinement. By providing explicit gradient information during the refinement process, AgentDrug enables targeted molecular modifications while maintaining similarity constraints. Our comprehensive experiments show that AgentDrug consistently outperforms powerful baselines on both single- and multi-property optimization tasks. It is particularly effective for complex multi-property objectives and can balance well with optimization accuracy and molecular similarity.
@@
Appendix A Evaluation Settings
[…]
@@ Appendix A Evaluation Settings
Box A1: Template for wrapping an objective in natural language to form the prompt Given [a given molecule $m$], modify it to [increase or decrease] its [desired properties $\{p_{i}\}_{i=1}^{N}$] by [thresholds $\{|d_{i}|\}_{i=1}^{N}$], respectively. Importantly, the modified molecule must be similar to the given one. Respond with only the SMILES string of the modified molecule. No explanation is needed.
Appendix B Ablation Studies
@@
ChatDrug AgentDrug ChatDrug AgentDrug
accuracy accuracy accuracy accuracy
@@
+LogP +TPSA l l 06.5 10.2 08.0 11.2
s s 01.6 07.4 05.6 09.6
+LogP -TPSA l l 07.2 11.6 11.8 09.6
s s 06.2 05.3 08.1 03.6
-LogP +TPSA l l 09.3 15.8 24.6 35.2
s s 03.8 08.4 13.3 21.8
-LogP -TPSA l l 06.3 05.0 08.0 02.9
s s 01.3 00.7 02.8 01.1
+LogP +QED l l 09.6 12.3 12.2 13.6
s s 02.2 01.6 01.3 01.8
+LogP -QED l l 13.3 23.3 21.9 27.7
s s 05.8 07.6 08.0 07.9
-LogP +QED l l 08.3 09.1 14.2 13.5
s s 01.3 00.7 01.4 01.5
-LogP -QED l l 11.2 09.0 25.0 31.2
s s 03.0 04.4 07.8 08.7
Average l l 09.0 12.0 15.7 18.1
s s 03.1 04.5 06.1 07.0
@@
B.1 Impact of using the example molecule
Table 5 presents results for both AgentDrug and ChatDrug when $m_{e}$ is excluded.
B.2 Impact of the number of iterations
Table 6 presents results for both methods using $T=4$, 55, and $6$ iterations, respectively.
Appendix C Ablation for the Database
Table 7: The results of AgentDrug with various sizes of the database D.
$p$ $d$ Qwen2.5-3B Qwen2.5-7B
5K 10K 20K 5K 10K 20K
accuracy accuracy accuracy accuracy accuracy accuracy
+LogP l 47.7 47.3 55.9 67.2 66.7 68.1
-LogP l 36.6 38.8 41.0 77.3 71.5 73.0
+TPSA l 41.8 41.2 43.6 64.5 62.8 65.0
-TPSA l 46.2 45.7 52.0 69.8 65.3 64.5
Average l 43.1 43.2 48.1 69.7 66.6 67.6
+LogP +TPSA l l 28.0 31.7 21.2 49.6 46.2 51.6
+LogP -TPSA l l 20.1 19.7 22.1 23.4 25.6 24.8
-LogP +TPSA l l 18.9 29.6 22.6 65.0 54.8 55.2
-LogP -TPSA l l 09.3 07.0 05.8 14.1 10.8 08.9
Average l l 19.1 22.0 17.9 38.0 34.4 35.1
In our main experiments, the database $\mathcal{D}$ consists of 10K molecules. Here, we conducted additional experiments with smaller (5K) and larger (20K) databases to thoroughly observe the behavior of AgentDrug. Table 7 displays the results of AgentDrug (with various sizes of the database $\mathcal{D}$) on LogP and TPSA using Qwen2.5 3B and 7B with loose thresholds.
We can observe that the size of the database $\mathcal{D}$ has a nonlinear impact on the performance of AgentDrug, and the impact depends on target properties. This might indicate that the property distributions of molecules in the database are more crucial than the size. Importantly, AgentDrug still performs well with a small database (5K), demonstrating its robustness.
Appendix D Related Work
Recent advancements in AI, particularly in large language models (LLMs), have begun to transform the drug discovery process by efficiently scaling many core molecular tasks Mak and Pichika (2019). Notably, Seidl et al. (2023); Zhao et al. (2023); Liu et al. (2024b) have incrementally pre-trained LLMs to advance molecular property prediction, synthesis, and retrosynthesis. In addition, Edwards et al. (2022); Liu et al. (2023b); Li et al. (2024); Le et al. (2024) have trained LLMs for molecular captioning. However, molecular optimization, which plays a critical role in designing improved compounds, remains underexplored.
Appendix E Discussion
In this section, we give a discussion on the motivation and nature of molecular optimization. Despite the existence of huge chemical databases like ZINC, ChEMBL, PubChem, or commercial catalogs containing billions of molecules, design and optimization are still essential for several reasons:
  • The estimated drug-like chemical space is > $10^{60}$ molecules. Even huge databases like ZINC, ChEMBL, and PubChem are just a small fraction of that space. Thus, the chance that the target molecule (satisfying multiple objectives) is already enumerated is low. Moreover, retrieval over huge molecule databases is extremely expensive because computing similarity between molecules is hard to parallelize.
  • Novelty for patentability. Pharmaceutical companies need novel chemical compounds they can patent. Retrieving a molecule already reported in the literature often means that they cannot claim exclusivity, making it commercially unviable.
In summary, retrieval from databases is only a valuable starting point to augment design and optimization. For example, AgentDrug leverages a small database (10K molecules) to augment the optimization process. It depends on the size of the database; the molecule $m_{e}$ can be easier or harder to find. Also, the retrieval may fail or return a low-similarity molecule in sparse areas.
Appendix F Full Algorithm
@@
Appendix G Examples
Illustrative examples of AgentDrug’s workflow, including detailed prompts, are provided in Figure 5 and Figure 6. The gradient signal used to guide the LLM is highlighted in red for clarity.